OpenStudy (anonymous):

Solve 3^ (1/2) * 3 ^ (1/2) !!!!

OpenStudy (anonymous):

@hartnn could you help me?

OpenStudy (anonymous):

Sorry, its Simplify, not Solve.

OpenStudy (anonymous):

PLZ HELLP ME

OpenStudy (anonymous):

What is thirty−six percent of fifty? one point four zero point seven two eighteen seventy−two

hartnn (hartnn):

firstly \(\large a \times a = a^2\) here, a= \(3^{1/2}\)

OpenStudy (anonymous):

I WILL GIVE U A MEDAL

OpenStudy (anonymous):

What is thirty−six percent of fifty? one point four zero point seven two eighteen seventy−two

OpenStudy (anonymous):

Recall that raising something to the power of a half is equivalent to square rooting. \[\sqrt3 \times \sqrt3\] Note anything interesting?

OpenStudy (anonymous):

actually the bases are same so the powers will be added because there is a multiply sign

hartnn (hartnn):

so, \( \large 3^{1/2} \times 3^{1/2} = 3^{2/2} = ...\)

OpenStudy (anonymous):

@hartnn okay, soo... would your answer just be 3.?

hartnn (hartnn):

YES! ")

OpenStudy (anonymous):

the answer is three with power 1

OpenStudy (anonymous):

Omg I feel so silly for not knowing! thanks everyone!!! :)

OpenStudy (anonymous):

ur welcome

hartnn (hartnn):

welcome ^_^

OpenStudy (anonymous):

is it indices???